# Ensemble Debiasing Across Class and Sample Levels for Fairer Prompting Accuracy

### v1

@@ 4.2 Overall Probability Correction Scheme
The overall probability correction scheme demonstrates the ratio of class-level corrections to sample-level corrections among all classes. Which type of correction should be used for a class is automatically determined by our method. Recall that:
Class-level correction: This correction involves adjusting the output class probability by multiplying it with a weight coefficient (determined through optimization). Probabilities of the same output class are corrected uniformly using the same weight across all samples. It operates at the class level, meaning it does not account for variations in initial probabilities between different samples.
Sample-level correction: This is a more refined correction utilizing membership functions to correct probabilities at a finer granularity. Instead of applying a single coefficient to each class probability, a unique mapping function is used for each class (determined through optimization). This allows for varying degrees of correction across different samples - some probabilities may be adjusted significantly, others minimally, and some may remain unchanged. This sample-level correction captures the nuances of individual samples, leading to more precise and tailored adjustments.
By combining these two correction approaches, our method ensures a balanced and adaptive refinement of class probabilities, enhancing the performance of the model.
@@
We exemplify the overall probability correction scheme with the $N$-shot case in Table 2. Numbers of classes are averaged over three runs of different demonstrations. As expected, most classes select class-level weight corrections. Membership corrections are less selected, but weakest classes benefit greatly from them, where 6 out of 7 tasks have selected membership corrections to correct their weakest classes (at least once over three runs). This demonstrates the necessity of enabling sample-level corrections to boost the performance of weak classes, and letting the optimization process decide if sample-level corrections are needed.
@@ 4.3 Larger Models on DDI (Biomedical-domain)
Due to effectively correcting probabilities for low-accuracy classes as shown by Table 3, our method demonstrates significant accuracy enhancements on DDI, across both small and large model variants.
More Llama-2-70B performance gains. Our method also brings significant performance gains to Llama-2-70B on the rest of the evaluation benchmarks. Detailed scores can be found in Appendix C.
@@ 5 Conclusion and Limitations
This paper proposes DCS, a post-hoc debiasing framework that unifies sample and class level corrections to enrich low-accuracy classes in ICL outputs. It achieves SoTA accuracy on seven benchmark text classification tasks, while effectively mitigating class performance imbalances. The resulting probability correction scheme clearly shows that combining membership correction and weight correction is essential to elevate weak classes. For limitations and future work, the post-hoc correction framework can be extended to other useful scenarios, such as text generation. Combining DCS into the decoding process can help rectify the probabilities of top candidates for the first token (or first several tokens) to lead a better, less biased response. Furthermore, this work deals with prediction biases when directly prompting LLMs. When LLMs or smaller LMs like BERT [30] are finetuned on tasks, which may result in further imbalanced class accuracies, post-hoc corrections can make class accuracies fairer. It would be valuable to understand how these output class probabilities differ from those by direct prompting, and enable custom extensions to the current framework to provide more targeted corrections. It would also be interesting to see how it may help with data-filtered or data-augmented finetuning. In terms of interpretability, those classes that apply class-level correction to any instance is interpretable only at the broader level, so further investigations on how to enforce more interpretability to weight corrections could be conducted.
Appendix A List of notations
@@
Appendix B Experimental Setups
@@
Appendix C Full Performance Comparisons on Llama-2-70B

### v2

@@ 4.2 Overall Probability Correction Scheme
The overall probability correction scheme demonstrates the ratio of class-level corrections to sample-level corrections among all classes. Which type of correction should be used for a class is automatically determined by our method. Recall that: Class-level correction: Probabilities of the same output class are corrected uniformly using a same correction weight across all samples. It operates at the class level, not accounting for sample-based differences. Sample-level correction: This is a more refined correction using membership correction functions. Instead of a single weight, a unique mapping function is applied for each class. This allows varying degrees of correction across different samples - some probabilities may be adjusted significantly, others minimally, and some may remain unchanged - capturing the nuances of individual samples, leading to more precise adjustments. By combining them, our method ensures a balanced and adaptive correction, enhancing performances.
@@
We exemplify the overall probability correction scheme with the $N$-shot case in Table 2. Numbers of classes are averaged over three runs of different demonstrations. As expected, most classes select class-level weight corrections. Membership corrections are less selected, but weakest classes benefit greatly from them, where 6 out of 7 tasks have selected membership corrections to correct their weakest classes (at least once over three runs). This demonstrates the necessity of enabling sample-level corrections to boost the performance of weak classes, and letting the optimization decide if sample-level corrections are needed.
@@ 4.3 Larger Models on DDI (Biomedical-domain)
Due to effectively correcting probabilities for low-accuracy classes as shown by Table 3, our method demonstrates significant accuracy enhancements on DDI, across both small and large model variants. For more Llama-2-70B performance gains, detailed scores can be found in Appendix D.
@@ 5 Conclusion and Limitations
This paper proposes DCS, a post-hoc debiasing framework that unifies sample and class level corrections to enrich low-accuracy classes in ICL outputs. It achieves SoTA accuracy on seven benchmark text classification tasks, while effectively mitigating class performance imbalances. The resulting probability correction scheme clearly shows that combining membership correction and weight correction is essential to elevate weak classes. For limitations and future work, the post-hoc correction framework can be extended to other useful scenarios, such as text generation. Combining DCS into the decoding process can help rectify the probabilities of top candidates for the first token (or first several tokens) to lead a better, less biased response. More future directions are discussed in Appendix E.
Appendix A Discussion on Solution Methods for NIP
Typical solutions for combinatorial optimization problems include the branch-and-bound (B&B) algorithm, which is useful for solving mixed-integer linear programs [48]. Solvers are made into both open-source software such as CVXPY [49, 50] or SCIP [51], and commercial solvers like Gurobi [52] or LINGO [53].
We need to point out that B&B and other related algorithms are not very suitable to search solutions for the proposed nonlinear integer programming based optimization model, which contain discontinuous functions and is inherently non-differentiable. Relaxation to subproblems is needed if B&B has to be used. Relaxation can be done by creating Lagrange multipliers and Lagrangian functions for each constraint, which is often easier to solve than the original nonlinear integer programming problem, because only continuous variables are involved [54, 55]. The relaxed problem finds the optimal value of the Lagrangian function, which is a lower bound for the original problem, and this optimal value of the relaxed problem is used as a bound in the B&B search. However, solving the subproblems efficiently becomes a challenge, as B&B is generally not a polynomial-time algorithm.
Instead, we can opt for metaheuristics [56], such as simulated annealing [35, 36], to find the optimal solution. Although heuristic search can get slower with increasing search space size [57], its computational complexity is within polynomial time [5]. To enable application of our proposed method to tasks with even more classes, we adopt simulated annealing (SA) to solve our model.
In SA, we define an initial temperature $T$ and an annealing schedule including the cooling rate and the stopping criteria. In each iteration, a neighborhood solution is generated by perturbing the current solution. The algorithm calculates the objective function value for the new solution (Step Evaluate Z in Fig. 2). If the new solution is better, accept it; if the new solution is worse, accept it with a probability determined by the Metropolis criterion. Then the temperature is reduced according to the cooling rate, until convergence. In addition, the initial solution can also be randomly generated, where $N$ numbers are randomly selected from $\mathbb{D}=\{1,\dots,D_{F}+D_{W}\}$ as the initial values for the selection variables $\xi_{j}$’s.
Appendix B List of notations
@@
Appendix C Experimental Setups
@@
Appendix D Full Performance Comparisons on Llama-2-70B
@@
Appendix E More Future Directions
This work deals with prediction biases when directly prompting LLMs. When LLMs or smaller LMs like BERT [58] are finetuned on tasks, which may result in further imbalanced class accuracies, post-hoc corrections can make class accuracies fairer. It would be valuable to understand how these output class probabilities differ from those by direct prompting, and enable custom extensions to the current framework to provide more targeted corrections. It would also be interesting to see how it may help with data-filtered or data-augmented finetuning. In terms of interpretability, those classes that apply class-level correction to any instance is interpretable only at the broader level, so further investigations on how to enforce more interpretability to weight corrections could be conducted.